# Looking To The Horizon: Probing Evolution in the Black Hole Spectrum With GW Catalogs

### v1

@@ IV Results
Figure 3: Rate density as a function of primary mass $m_{1}$ and luminosity distance $d_{L}$, assuming a power-law distribution for $m_{2}$. Red dots represent the medians of parameter estimation (PE) samples for BBH events, while the orange contours correspond to $p_{\mathrm{det}}$ levels. (One event, GW190805_211137, has a median $d_{L}\sim 7500$ Mpc outside the range of the plot.)
@@
To clearly display any evolution over redshift, we extract the rate density over $m_{1}$ at a sequence of $d_{L}$ or redshift values.
Figure 4: Merger rate density (median estimate) as a function of the primary mass $m_{1}$, with different curves representing redshift values indicated by color, assuming a power law distribution of mass ratio. Estimates at low mass are increasingly affected by selection bias towards high redshift.
We show median rate values in Fig. 4, and corresponding 90% uncertainty bands (applying cuts on large uncertainties) in Fig. 5.
Figure 5: Merger rate density as a function of primary mass with symmetric 90% confidence intervals, corresponding to the results shown in Fig. 4. Each curve shows a constant distance or redshift, shown by the color scale, with constant offsets applied for clarity. We only plot values where both (1) the 5th percentile is $>10^{-3}\times$ median, and (2) the 95th percentile is 200$\leq$ 200$\times$ 200 × the median.
We observe two main apparent trends in Fig. 4: first, the height of the $10\,M_{\odot}$ peak, and the distribution below $30\,M_{\odot}$ as a whole, decreases rapidly at higher redshift; however this can be entirely attributed to “censorship” in this region due to the selection function. Second, the falloff at high mass ($>\!50\,M_{\odot}$) becomes shallower at higher redshift. Although this region is fully accessible to the detector network, Fig. 5 indicates that rate uncertainties are extremely large, which is not surprising given the very low statistics of detected high-mass events. Hence overall, this analysis is consistent with no evolution of the primary mass distribution.
@@
Figure 6: BBH rate density as a function of component mass $m_{1}$ and luminosity distance $d_{L}$, for fixed $m_{2}$. Note that, for this plot only, we do not impose $m_{1}>m_{2}$ in displaying the KDE. The orange contours represent $p_{\mathrm{det}}$ at the given $m_{2}$ values. Red plus markers indicate the median of PE samples from GWTC-3.
From the 3D results, we compute rates for fixed values of $m_{2}$, as shown in Fig. 6 for $m_{2}=(10,35)\,M_{\odot}$, overplotting contours of $p_{\mathrm{det}}$. (Here, unlike for the rest of our results, we do not impose $m_{2}<m_{1}$: thus, these plots show a slice through the full 2-d component mass plane.) For $m_{2}=10\,M_{\odot}$, the overdensity around $m_{1}\simeq 10\,M_{\odot}$ persists up to the limit of detected events around 111\,1Gpc ($z\simeq 0.2$). For $m_{2}=35\,M_{\odot}$, the clear peak at $m_{1}\simeq 35\,M_{\odot}$ persists out to a distance $\sim 4\,$Gpc ($z\simeq 0.6$), however detected events fall somewhat short of the $p_{\mathrm{det}}=0.1$ contour, indicating that caution should be used when interpreting the reconstruction at such large distances. These features are roughly as expected from the 2D results shown in Fig. 3. Our comoving rate estimates also show an apparent increase at very low distance $d_{L}\lesssim 0.3\,$Gpc: this is an artefact of our choice to construct a KDE over $d_{L}$, as the kernels have nonzero support at zero distance. Hence, we do not expect the estimate to be valid at distances below the closest detected events ($\sim\!0.3\,$Gpc).
Figure 7: Rate density as a function of component masses at fixed luminosity distance or redshift values. The red dots represent the medians of parameter estimation (PE) samples.
We also select different $d_{L}$ values to plot the median rate over component masses, as in Fig. 7. The distribution at 0.90.90.9\,0.9Gpc shown in the upper plot is essentially unchanging for distances out to the detection horizon for $10\,M_{\odot}$ components; the lower plot suggests a shift towards apparent higher support for high BH masses at larger distances, however at 1.81.81.8\,1.8Gpc the principal change is again a suppression of the low-mass component due to selection effects (censorship).
Figure 8: Merger rate density (median estimate) as a function of the primary mass $m_{1}$, integrated over secondary mass, with different curves representing rates at redshift values indicated by color.
We further investigate potential evolution by taking several fixed $d_{L}$ (or $z$) values and integrating the joint rate over $m_{2}$ to obtain rate vs. primary mass $m_{1}$. Median rates are shown in Fig. 8 and rates with uncertainties in Fig. 9, showing similar trends to the 2D case where a power-law assumption was applied to $m_{2}$. Fig. 8 suggests a more visible trend towards shallower fall-off of the high mass distribution with increasing redshift; however, uncertainties remain extremely large for $m_{1}>50\,M_{\odot}$. Hence, again we do not see positive evidence for evolution.
@@
One further common feature of our rate estimates is that they slightly decrease with redshift, even well within the “horizon” of detected events where $p_{\mathrm{det}}\gtrsim 0.1$, except for the visible increase at high masses. This is in apparent contrast to some models of redshift evolution, for instance taking the comoving rate to vary proportional to $\mathcal{R}_{0}(1+z)^{\kappa}$ a generally positive, though small value of $\kappa$ is inferred Abbott et al. (2023b). However, non-parametric approaches to measuring evolution of the total BBH merger rate do not confirm a clear increase up to small redshifts $\lesssim 0.5$, some even showing a slight decrease Edelman et al. (2023); Callister and Farr (2024); Ray et al. (2023). Only for $z\gtrsim 0.6$ is an increasing trend visible: but as seen e.g. in Fig. 3, high-redshift estimates are driven almost entirely by high-mass events.
Such apparent discrepancies may require further investigation, as finite, though hopefully small, biases can arise, either in our method or in (semi-)parameterized models. On the KDE side the choice of a constant kernel (up to local adaptive scaling) in the logarithm of masses and in luminosity distance may cause some bias given the current low number of detections; we expect such biases to be reduced with higher event counts, as optimal bandwidths will become smaller and the exact choice of kernel becomes less important. Conversely parameterized evolution models, or even those allowing for more general $z$-dependence, typically assume specific functional forms for many parameters such as mass ratio and spins. Even when dependence on every parameter is described non-parametrically Edelman et al. (2023), any assumption that different parameters are uncorrelated is a significant restriction. Such assumptions may lead to biases in estimating the selection function or in individual event parameters, which may partially mimic an evolving rate, since rate estimates at different redshifts are informed by events with different intrinsic properties.
@@ Appendix C Conversion to Cosmological Rate Density
$\mathrm{d}\mathcal{R}=\frac{\mathrm{d}N}{\mathrm{d}d_{L}\mathrm{d}t_{\mathrm{det}}}\frac{\mathrm{d}t_{\mathrm{det}}}{\mathrm{d}t_{\mathrm{s}}}\frac{\mathrm{d}d_{L}}{\mathrm{d}z}\left(\frac{\mathrm{d}V_{c}}{\mathrm{d}z}\right)^{-1}.\vspace*{0.05cm}$ (4)
We have $\mathrm{d}t_{\mathrm{det}}/\mathrm{d}t_{\mathrm{s}}=1+z$ due to time dilation. Further, for flat $\Lambda$CDM cosmology we have (e.g. Hogg (1999))

### v2

@@ IV Results
Figure 3: Rate density as a function of primary mass $m_{1}$ and luminosity distance $d_{L}$, assuming a power-law distribution for $m_{2}$. Red ++ symbols represent the medians of PE samples for GWTC-3 BBH events, while the orange contours correspond to $p_{\mathrm{det}}$ levels. One event, GW190805_211137, has a median $d_{L}\sim 7500$ Mpc outside the range of the plot.
@@
To display any evolution over redshift, we extract the rate density over $m_{1}$ at a sequence of $d_{L}$ or redshift values.
Figure 4: Top panel: Merger rate density (median estimate) as a function of the primary mass $m_{1}$, with different curves representing redshift values indicated by color, assuming a power law distribution of mass ratio. Estimates at low mass are increasingly affected by censorship towards high redshift: corresponding regions with $p_{\mathrm{det}}<0.1$ are shown as dotted curves. Bottom panel: Rate density for a subset of redshift values, with shaded 90% uncertainty intervals. We omit values with excessively high uncertainty for clarity (as also in Fig. 5).
We show median rate values in Fig. 4, with 90% uncertainty bands for selected redshifts (omitting points with extremely large uncertainty), and show uncertainties for all redshifts in Fig. 5.
Figure 5: Merger rate density as a function of primary mass with symmetric 90% confidence intervals, corresponding to the results shown in Fig. 4. Each curve shows a constant distance or redshift, shown by the color scale, with constant offsets applied for clarity. We only plot values where both (1) the 5th percentile is $10^{-3}\times$ median, and (2) the 95th percentile is $\leq 200\times$ the median.
We observe two main apparent trends in Fig. 4: first, the height of the $10\,M_{\odot}$ peak, and the distribution below $30\,M_{\odot}$ as a whole, decreases rapidly at higher redshift; however this can be largely attributed to “censorship” in this region, i.e. absence of detected events as expected for very small detection probability. At intermediate masses ($m_{1}\sim 35\,M_{\odot}$) there is a weak apparent trend toward decreasing rate with redshift. However the uncertainty bands in the lower plot still overlap for mass ranges unaffected by censorship, hence we do not find the trend to be significant. Second, the falloff at high mass ($50\,M_{\odot}$) becomes shallower at higher redshift. Although this region is fully accessible to the detector network, Fig. 5 indicates that rate uncertainties are very large, as expected given the low statistics of detected high-mass events. Overall, this analysis is consistent with no evolution of the primary mass distribution given uncertainties due to finite statistics.
@@
Figure 6: BBH rate density as a function of component mass $m_{1}$ and luminosity distance $d_{L}$, for fixed $m_{2}$. Note that, for this plot only, we do not impose $m_{1}>m_{2}$ in displaying the KDE. The orange contours represent $p_{\mathrm{det}}$ at the given $m_{2}$ values. Red ++ markers indicate medians of PE samples.
From the 3D results, we compute rates for fixed values of $m_{2}$, as shown in Fig. 6 for $m_{2}=(10,35)\,M_{\odot}$, overplotting contours of $p_{\mathrm{det}}$. (Here, unlike the rest of our results, we do not impose $m_{2}<m_{1}$: thus these plots show a slice through the full 2-d component mass plane.) For $m_{2}=10\,M_{\odot}$, the overdensity around $m_{1}\simeq 10\,M_{\odot}$ persists up to the limit of detected events around 11\,Gpc ($z\simeq 0.2$). For $m_{2}=35\,M_{\odot}$, the clear peak at $m_{1}\simeq 35\,M_{\odot}$ persists out to a distance $\sim 4\,$Gpc ($z\simeq 0.6$), however detected events fall somewhat short of the $p_{\mathrm{det}}=0.1$ contour, indicating that caution should be used when interpreting the reconstruction at such large distances. These features are roughly as expected from the 2D results shown in Fig. 3. Our comoving rate estimates also show an apparent increase at very low distance $d_{L}\lesssim 0.3\,$Gpc: this is an artefact of our choice to construct a KDE over $d_{L}$, as the kernels have nonzero support at zero distance. Hence, we do not expect the estimate to be valid at distances below the closest detected events ($\sim\!0.3\,$Gpc).
Figure 7: Rate density as a function of component masses at fixed luminosity distance or redshift values. Red ++ markers represent the medians of PE samples.
We also select different $d_{L}$ values to plot the median rate over component masses, as in Fig. 7 (note the different color scales between the two panels). The distribution at 0.90.9\,Gpc shown in the upper plot is essentially unchanging for distances out to the detection horizon for $10\,M_{\odot}$ components; the lower plot suggests a shift towards apparent higher support for high BH masses at larger distances, however the principal change here is again a suppression of the low-mass component due to very small $p_{\mathrm{det}}$ (censorship).
Figure 8: Merger rate density (median estimate) as a function of the primary mass $m_{1}$, integrated over secondary mass, with different curves representing rates at redshift values indicated by color. Regions with $p_{\mathrm{det}}<0.1$, estimated as in Fig. 4, are shown as dotted curves.
We further investigate potential evolution by taking several fixed $d_{L}$ (or $z$) values and integrating the joint rate over $m_{2}$ to obtain rate vs. primary mass $m_{1}$. Median rates are shown in Fig. 8 and rates with uncertainties in Fig. 9, showing similar trends to the 2D case where a power-law assumption was applied to $m_{2}$.
@@
Fig. 8 suggests a more visible trend towards shallower fall-off of the high mass distribution with increasing redshift; however, uncertainties remain extremely large for $m_{1}>50\,M_{\odot}$. Hence, again we do not see positive evidence for evolution.
While this conclusion agrees with most other studies, it is in contrast to Rinaldi et al. (2024) which claims non-trivial evolution, specifically decreasing low-mass rates, resp. increasing higher-mass rates with redshift. While the present work shares a method of estimating selection effects with Rinaldi et al. (2024), i.e. the injection fit of Lorenzo-Medina and Dent (2025), our inference on the detected BBH distribution is obtained with a completely different and independent analysis. We observe similar trends, however the large uncertainties we obtain over much of the parameter space imply the trends are not significant.
@@
We also conduct an verification analysis imposing a stricter bound on FAR of 0.25/0.25/y to check for possible bias due to noise contamination of our event set  [^2] We thank the referee for suggesting this check.: as shown in Appendix D, we find no significant changes in the overall inference, although statistical uncertainties increase due to the lower detection count.
One further common feature of our rate estimates is that they slightly decrease with redshift, even well within the “horizon” of detected events where $p_{\mathrm{det}}\gtrsim 0.1$, except for a possible increasing trend up to $z\sim 0.5$ at high masses. This is in apparent contrast to some models of redshift evolution, for instance, taking the comoving rate to vary proportional to $\mathcal{R}_{0}(1+z)^{\kappa}$ a generally positive, though small value of $\kappa$ is inferred Abbott et al. (2023b). However, non-parametric approaches to measuring evolution of the total BBH merger rate do not confirm an increasing rate at small redshifts $\lesssim 0.5$, some even showing a slight decrease Edelman et al. (2023); Callister and Farr (2024); Ray et al. (2023). Only for $z\gtrsim 0.6$ is an increasing trend visible in such studies – though not in our results. As seen e.g. in Fig. 3, high-redshift estimates must be driven almost entirely by high-mass events.
Such apparent discrepancies may require further investigation, as finite, though hopefully small, biases can arise, either in our method or in (semi-)parameterized models. On the KDE side the choice of a constant kernel (up to local adaptive scaling) in the logarithm of masses and in luminosity distance may cause some bias given the current low number of detections; we expect such biases to be reduced with higher event counts, as optimal bandwidths will become smaller and the exact choice of kernel becomes less important. Conversely, parameterized evolution models, or even those allowing for more general $z$-dependence, typically assume specific functional forms for many parameters such as mass ratio and spins. Even when dependence on every parameter is described non-parametrically Edelman et al. (2023), any assumption that different parameters are uncorrelated is a significant restriction. Such assumptions may lead to biases in estimating the selection function or in individual event parameters, which may partially mimic an evolving rate, since rate estimates at different redshifts are informed by events with different intrinsic properties. Our current analysis also does not allow for the BBH effective spin distribution to evolve over redshift; nontrivial evolution, as suggested in Biscoveanu et al. (2022), could impact inferences at high $z$ via the selection function.
@@ Acknowledgements
DATA AVAILABILITY
The data that support the findings of this article are openly available at LVK (2021a) and associated analysis data are available at https://zenodo.org/records/17178478.
@@ Appendix C Conversion to Cosmological Rate Density
$\mathrm{d}\mathcal{R}=\frac{\mathrm{d}N}{\mathrm{d}d_{L}\mathrm{d}t_{\mathrm{det}}}\frac{\mathrm{d}t_{\mathrm{det}}}{\mathrm{d}t_{\mathrm{s}}}\frac{\mathrm{d}d_{L}}{\mathrm{d}z}\left(\frac{\mathrm{d}V_{c}}{\mathrm{d}z}\right)^{-1}.\vskip 1.42271pt$ (5)
[…]
@@
Appendix D Analysis of Events With FAR $\leq 0.25/$y
Motivated by the possibility of noise contamination in our event set, we apply a stricter cut on the FAR in the selection of BBH events to examine its impact on the results. With this threshold set to $0.25$ per year, we obtain a total of 62 BBH events from the GWTC-3 catalog (we also recalculate $p_{\mathrm{det}}$ corresponding to this more stringent threshold). Here we present results derived from these selected events.
As shown in Figure 15, we obtain similar results to our main analysis. However, we see much larger uncertainties for $m_{1}\gtrsim 15\,M_{\odot}$ at low redshifts. Such differences may be driven by a preference for different KDE bandwidth and adaptive parameter values as compared to our main analysis, given a smaller number of detections occupying a narrower overall range of $d_{L}$.
Figure 15: Merger rate density as a function of the primary mass $m_{1}$ with uncertainties, where only events with FAR $\leq$ 0.25 are included (62 BBH events).